Swagelok Enterprises is a manufacturer of miniature fittings and valves. Over a 5-year period, the costs associated with onei tried 5% and 25% it is wrong? should i use interpolation? if so, how?
Answer #1

First Cost = 22,000

Annual Cost = 18,000

Annual Revenue = 27,000

Net Annual Revenue = 27,000 - 18,000 = 9,000

Salvage Value = 5,000

Life = 5 years

Calculating rate of return using Trial and error Method

Let the interest rate = 32%. Calculate PW at 32%.

PW = -22,000+9,000 (P/A, 32%, 5) + 5,000 (P/F, 32%, 5)

PW = -22,000+9,000 (2.3452) + 5,000 (0.2495) = 354

The PW is positive. Increase the interest rate to 35% to get negative PW.

PW = -22,000+9,000 (P/A, 35%, 5) + 5,000 (P/F, 35%, 5)

PW at 35% = -22,000+9,000 (2.2200) + 5,000 (0.2230) =-905

Using interpolation

ROR = 32% + [354 – 0 ÷ 354 – (-905)]*3% = 32.84%

The rate of return that the company made on the product is 32.84%